OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality -2x - y >= -8 Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (texaschic101):

-2x - y >= -8 -y >= 2x - 8 -- multiply by -1 to make y positive, and change the sign y <= -2x + 8 If it contains an equal sign, it will be a solid line. And it will be shaded below the line because of the < sign. y <= -2x + 8 (sub in points (0,0) 0 <= -2(0) + 8 0 <= 0 + 8 0 <= 8 (correct) so it does contain the origin answer is : D

OpenStudy (anonymous):

thank you again!!